What is the intersection of these two lines?
2x+4=y and x-y=2
1
Expert's answer
2012-09-14T10:38:07-0400
We need to solve a system of equations:

2x+4 = y (1)

x - y = 2 (2)

Let's substituted y from (1) into the second equation:

x - (2x + 4) = 2

-x - 4 = 2

x = -6

And then count y: 2*(-6) +4 = -8
It's a point with coordinates (-2;-8).